OpenStudy (angel_kitty12):

I'm having a problem with probability. Please someone help ASAP. A first grade class of nine girls and seven boys walks into a class in alphabetical order (by last name). What is the probability that three girls are the first to enter the room? Show your calculation. (A) 0.15 (B) 0.20 (C) 0.35 (D) 0.45

OpenStudy (jack1):

9girls + 7boys= 16 students calc probability 1st person is a girl and 2nd person is a girl and 3rd person is a girl...

OpenStudy (angel_kitty12):

So 9/16 • 8/16 • 7/16 ?

OpenStudy (jack1):

so...P(g AND g and g) = ? P(1st is a girl) = 9 girls/ 16 students =9/16 P(2nd is a girl) = 8girls (remaining)/15 students (remaining outside) = 8/15 now u : P(3rd is a girl) = ?? finally: (u again) P(1st is g AND 2nd is g and 3rd is g) = P(1st is g x 2nd is g x 3rd is g) = =???

OpenStudy (angel_kitty12):

So it's 9/16•8/15•7/14 which will be 0.15

OpenStudy (jack1):

yep, perfect!